OpenStudy (brycehale):

ABCD is a parallelogram. If AB = 2x - y, CD = 7, BC = x + y , AD = 2 find x and y.

OpenStudy (brycehale):

I have gotten x = 9 and y = -7

OpenStudy (brycehale):

But when I checked it is was wrong

OpenStudy (anonymous):

In a parallelogram, opposite sides are equal. So AB and CD are equal, AD and BC are equal. 2x - y = 7 x + y = 2 Solving the equation by adding them you get 3x = 9 or x = 3 And replacing x value in one of the equations you get 3 + y = 2 or y = -1 Is x = 3 and y = -1 correct?

OpenStudy (brycehale):

It was! Thanks! I was over-calulating!

OpenStudy (anonymous):

No problem (: haha good luck buddy

OpenStudy (brycehale):

Thanks. And I was using the substitution method instead of the addition method as you used.